OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half.

OpenStudy (anonymous):

a. y - 4 = one half(x - 1) b. y + 4 = 2(x - 1) c. y + 4 = one half(x - 1) d. y - 4 = 2(x + 1)

OpenStudy (anonymous):

a or c rigjt?

OpenStudy (anonymous):

Use the point slope formula do you know what it is and how to use it?

OpenStudy (anonymous):

y1-y2/x1-x2

OpenStudy (anonymous):

That is the formula to find the slope of a line. Once you have the slope and a point you can use the point slope formula to find the rest of the equation. Those answers are all in point slope form. Give me a moment to work it out and ill get back too you

OpenStudy (anonymous):

If you use the point slope m=y-y coordinate/x-x coordinate then you get 1/2=y-(-4)/x-1... Multiply both sides by (x-1) and you get 1/2(x-1)=y+4 or c make sense?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

confused lol

OpenStudy (anonymous):

m stands in for the slope try to write it out and replace my / with fraction bars and it should be a little easier to understand

OpenStudy (abdullahm):

yes, you have correctly narrowed it down to a or c. They gave you a point \(\sf (1,-4)\), and that is in the form of \(\sf (x_1,y_1)\) So, plug in the values into: \(\sf y-y_1=m(x-x_1)\) where m is the slope

OpenStudy (anonymous):

showing me up lol with all your fancy techniques

OpenStudy (anonymous):

so y+4= 1/2(x - 1)

OpenStudy (abdullahm):

correct

OpenStudy (abdullahm):

This is not fancy, this is the easy way @3714

OpenStudy (anonymous):

thanks again abdullah

OpenStudy (abdullahm):

np :)

OpenStudy (anonymous):

lol I tired

OpenStudy (abdullahm):

but your way would lead to the correct answer too, so no worries xD